Quiz 2: 1.) Write as a single simplified logarithm:

ln(1)ln(3/5) + 2ln(3) − ln(e)ln(4) =

2.) Graph the following

1a) y = e^{x}  (b) y = e^{−x}  (c) y = ln(x)

3.) Give an example of an initial value problem with multiple solutions.

4.) Define: A function f is linear if